Harry is 1/3 old as Ron and eight years younger than Hermoine. FInd the some of their Ages of Harry is 8 years old.
Answer by satyareddy22(84)

Let x,y,z be the ages of harry,ron,hermoine respectively.
x=y/3---------(1)
z-x=8---------(2)
Harry is 8 years old=>x=8yrs
from (1) 8=y/3=>y=8*3=24=>y=24yrs
z-8=8=>z=8+8=16=>z=16yrs
sum of their Ages=x+y+z=8+16+24=48yrs